# Learning Cache Coherence Traffic for NoC Routing Design

### v1

@@ Abstract.
The rapid growth of multi-core systems highlights the need for efficient Network-on-Chip (NoC) design to ensure seamless communication. Cache coherence, essential for data consistency, substantially reduces task computation time by enabling data sharing among caches. As a result, routing serves two roles: facilitating data sharing (influenced by topology) and managing NoC-level communication. However, cache coherence is often overlooked in routing, causing mismatches between design expectations and evaluation outcomes. Two main challenges are the lack of specialized tools to assess cache coherence’s impact and the neglect of topology selection in routing. In this work, we propose a cache coherence–aware routing approach with integrated topology selection, guided by our Cache Coherence Traffic Analyzer (CCTA). Our method achieves up to achieves up to 10.52% lower packet latency, 55.51% faster execution time, and 49.02% total energy savings, underscoring the critical role of cache coherence in NoC design and enabling effective co-design.
@@ 4. Experiment
The experiment, using the PARSEC 2.1 benchmark suite (b27, ) in the Gem5 simulator, evaluates performance in real-world scenarios. Energy consumption is measured with McPAT (b28, ), and simulations are conducted on NoC configurations with 16 and 64 cores. Table 1 provides a comprehensive overview of the system configuration.
All CCTA traces obtained from Gem5’s output file serve as environment states. Training is organized into multiple episodes, each representing a complete application run. At every step, the current state is fed into the Q-network to compute Q-values for candidate topologies. An $\epsilon$-greedy policy selects a topology, and the Link Weight Update module (a three-layer MLP) refines routing. The environment simulates the chosen configuration, updates NoC and cache coherence metrics, and computes a multi-objective reward that drives backpropagation in both networks, with the state updated before the next step. Although experiments focus on one application at a time, the framework can handle multiple workloads by randomly selecting benchmarks per episode while leaving the rest of the training process unchanged.
We compare methods that account for cache coherence with those that do not. For the latter, we evaluate dimension-order routing (XY), Q-learning-based methods like BiLCQ (b31, ) and CrQ (b32, ), as well as the congestion-adaptive DyAD (b33, ). While BiLCQ and CrQ use Q-tables for updates, DyAD makes decisions based on queue lengths without Q-learning. To ensure a fair comparison, we integrate our Cache Coherence Traffic Analyzer (CCTA) with RL-based (b29, ) and DRL-based (b30, ) methods(i.e., non-cc-aware RL and non-cc-aware DRL) transforming them into cache coherence-aware versions, referring to as cc-aware RL and cc-aware DRL. This enables direct comparison within a cache-coherent NoC environment. CCTA captures detailed coherence metrics, ensuring a thorough evaluation of each method’s traffic management, coherence handling, and overall efficiency.
@@ 4.2. Analysis of Cache Coherence Impact
Results on Total Energy. Figure 8(b) shows that our method lowers energy consumption by 44.75%, 48.26%, 49.02%, 47.38%, 41.77%, and 40.67% compared to XY, DyAD, BiLCQ, CrQ, cc-aware RL, and cc-aware DRL. Non-cc-aware methods consistently consume more energy across all workloads. Under cache coherence (cc-aware RL and DRL), all workloads except Vips match our energy and latency by effectively managing congestion and coherence traffic. However, they rely on less flexible topologies and incremental updates, limiting synergy between topology selection, routing, and cache coherence. Our method dynamically integrates these elements, avoiding unnecessary overhead and achieving notably shorter execution times (Figure 8(c)).
@@
Results on Execution Time. Our method (Figure 8(c)) reduces execution time by an average of 55.51% and 31.20% compared to cc-aware RL-based and cc-aware DRL-based methods, respectively. Notably, both cc-aware RL and DRL show nearly the same execution time for X264, Fer, and Bla. In these workloads, the communication phases tend to be more predictable or follow well-defined loops, enabling both methods to maintain comparable overhead and performance. By contrast, Flu and Vips exhibit more dynamic, bursty data-sharing patterns, leading the two cc-aware approaches to diverge in traffic management and produce greater variation in execution times.
Results on Cache Coherence Performance. To evaluate cache coherence performance, we measure cache coherence-driven message transmissions. As shown in Figure 9, our method reduces transmissions by 8.33% and 0.95% compared to cc-aware RL and DRL, respectively, while maintaining stability across all applications. Although cache coherence naturally increases traffic, our proactive strategy emphasizes congestion reduction and communication efficiency. For workloads like Fer and Bla, this approach slightly increases coherence messages to ensure stable data exchange and balanced packet distribution. Coupled with improved performance (Figure 8), our method delivers a superior trade-off among NoC efficiency, coherence traffic, and overall system performance. Moreover, RL- and DRL-based routing show similar performance under cache coherence for certain workloads, underscoring the need for more advanced cc-aware frameworks—such as ours—to handle cache coherence more effectively.
@@ 5. Conclusion
In this work, we demonstrate that cache coherence significantly reduces task computation time by enabling efficient data sharing among caches, improving overall system performance. Building on this insight, we address three key shortcomings in NoC design: (i) misalignment between design objectives and actual performance when considering cache coherence, (ii) lack of analysis tools, limiting the understanding of cache coherence’s impact, and (iii) unexploited potential of integrating topology selection into routing. To address these challenges, we propose a DRL-based framework that jointly optimizes topology selection and routing, integrating path learning with topology decisions while accounting for NoC and cache coherence metrics. Additionally, we introduce a Gem5-compatible analysis tool for precise cache coherence evaluation, providing deeper insights into its role in NoC design. Our approach achieves up to 10.52% lower packet latency, 55.51% faster execution time, and 49.02% total energy savings, while maintaining minimal coherence overhead. This highlights the importance of explicitly considering cache coherence in NoC design, paving the way for NoC–coherence co-design and demonstrating the broad potential of our methodology.

### v2

@@ Abstract.
The rapid growth of multi-core systems highlights the need for efficient Network-on-Chip (NoC) design to ensure seamless communication. Cache coherence, essential for data consistency, substantially reduces task computation time by enabling data sharing among caches. As a result, routing serves two roles: facilitating data sharing (influenced by topology) and managing NoC-level communication. However, cache coherence is often overlooked in routing, causing mismatches between design expectations and evaluation outcomes. Two main challenges are the lack of specialized tools to assess cache coherence’s impact and the neglect of topology selection in routing. In this work, we propose a cache coherence–aware routing approach with integrated topology selection, guided by our Cache Coherence Traffic Analyzer (CCTA). Our method achieves up to 10.52% lower packet latency, 55.51% faster execution time, and 49.02% total energy savings, underscoring the critical role of cache coherence in NoC design and enabling effective co-design.
@@
copyright: acmlicensedjournalyear: 2025copyright: ccconference: Great Lakes Symposium on VLSI 2025; June 30-July 2, 2025; New Orleans, LA, USAbooktitle: Great Lakes Symposium on VLSI 2025 (GLSVLSI ’25), June 30-July 2, 2025, New Orleans, LA, USAdoi: 10.1145/3716368.3735166isbn: 979-8-4007-1496-2/2025/06ccs: Networks Network on chip
@@ 4. Experiment
The experiment, using the PARSEC 2.1 benchmark suite (b27, ) in the Gem5 simulator, evaluates performance in real-world scenarios. Energy consumption is measured with McPAT (b28, ). Simulations are conducted on NoC configurations with 16 and 64 cores, as shown in Table 1. The system employs a directory-based MESI cache coherence protocol, which is both well-studied in academia and aligned with industry standards such as the Coherent Hub Interface (CHI), ensuring realistic modeling of coherence interactions.
@@
All CCTA traces extracted from Gem5’s output files are utilized as the state representations for the learning environment. The training process is structured into multiple episodes, where each episode corresponds to the execution of a complete application workload. At each decision step within an episode, the current state is passed to the Q-network, which evaluates the Q-values for candidate network topologies—including Crossbar, Mesh, Pt2Pt, Torus, Fat Tree, and Flattened Butterfly. An $\epsilon$-greedy policy then selects one of these topologies and the Link Weight Update module (a three-layer MLP) refines routing. The updated configuration is simulated in the Gem5 environment, which produces updated metrics for NoC traffic and cache coherence behavior. Based on these results, a multi-objective reward is computed that drives backpropagation for both the Q-network and the MLP-based refinement module, with the state updated before the next step. Although experiments focus on one application at a time, the framework can handle multiple workloads by randomly selecting benchmarks per episode while leaving the rest of the training process unchanged.
We compare methods that account for cache coherence with those that do not. For the latter, we evaluate dimension-order routing (XY), Q-learning-based methods like BiLCQ (b31, ) and CrQ (b32, ), as well as the congestion-adaptive DyAD (b33, ). While BiLCQ and CrQ use Q-tables for updates, DyAD makes decisions based on queue lengths without Q-learning. To ensure a fair comparison, we integrate our Cache Coherence Traffic Analyzer (CCTA) with RL-based (b29, ) and DRL-based (b30, ) methods(i.e., non-cc-aware RL and non-cc-aware DRL) transforming them into cache coherence-aware versions, referring to as cc-aware RL and cc-aware DRL. This enables direct comparison within a cache-coherent environment.
@@ 4.2. Analysis of Cache Coherence Impact
Results on Total Energy. Figure 8(b) shows our method reduces energy consumption by 44.75%, 48.26%, 49.02%, 47.38%, 41.77%, and 40.67% relative to XY, DyAD, BiLCQ, CrQ, cc-aware RL, and cc-aware DRL. Non-cc-aware techniques consistently consume more energy. Under cache coherence, cc-aware RL and DRL match our energy and latency on all workloads except Vips by effectively managing congestion and coherence traffic. However, they rely on inflexible topologies and incremental updates, limiting synergy between topology selection, routing, and coherence. Our method dynamically integrates these elements, avoiding unnecessary overhead and achieving notably shorter execution time (Figure 8(c)).
Results on Execution Time. Our method (Figure 8(c)) reduces execution time by 55.51% and 31.20% compared to cc-aware RL and cc-aware DRL, respectively. Notably, cc-aware RL and DRL achieve nearly identical execution times on X264, Fer, and Bla—workloads whose predictable, loop-driven communication phases allow both methods to incur comparable overhead and performance. In contrast, Flu and Vips exhibit more bursty, dynamic data-sharing patterns, causing the two methods to diverge in traffic management and produce greater variation in execution times.
Results on Cache Coherence Performance. Figure 9 shows our method reduces coherence-induced message transmissions by 8.33% and 0.95% compared to cc-aware RL and DRL, respectively, while remaining stable across all workloads. Although coherence protocols inherently increase traffic, our proactive approach throttles congestion and boosts communication efficiency.For workloads like Fer and Bla, this approach slightly increases coherence messages to ensure stable data exchange and balanced packet distribution. Coupled with improved performance (Figure 8), our method delivers a superior balance of NoC efficiency, coherence traffic, and overall system performance. Moreover, RL- and DRL-based routing show similar performance under cache coherence for certain workloads, underscoring the need for more advanced cc-aware frameworks like ours to handle coherence more effectively.
@@ 5. Conclusion
In this work, we demonstrate that cache coherence significantly reduces task computation time by enabling efficient data sharing among caches, thereby enhancing overall system performance. Building on this insight, we identify and address three critical gaps in current NoC design: (i) misalignment between design objectives and actual performance when considering cache coherence, (ii) lack of analysis tools, limiting the understanding of cache coherence’s impact, and (iii) unexploited potential of integrating topology selection into routing decisions. To address these challenges, we propose a DRL-based framework that jointly optimizes topology selection and routing by dynamically learning from both NoC and cache coherence metrics. Additionally, we develop a Gem5-compatible Cache Coherence Traffic Analyzer (CCTA) to enable precise, fine-grained evaluation of coherence behaviors, providing deeper insights into its role in NoC design. Our approach achieves up to 10.52% lower packet latency, 55.51% faster execution time, and 49.02% total energy savings, while maintaining minimal coherence overhead. These results highlight the importance of explicitly considering cache coherence in NoC design, paving the way for NoC–coherence co-design and demonstrating the broad potential of our methodology.
Acknowledgements.
This work is partially supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (RG94/23).